# Singleton-Optimized Conformal Prediction

### v1

@@ 1 Introduction
However, average size is not necessarily the ideal measure of efficiency. Often, the most desirable outcome is an unambiguous prediction, a singleton set containing only one label. A set of size two or more may require additional human intervention or changing the workflow when used in downstream analysis, and thus brings an outsized cost. This motivates an alternative efficiency criterion, first conceptualized in Vovk et al. (2005) as the M-criterion, which seeks to minimize the probability of producing a non-singleton set, $\mathbb{P}_{X}[|C(X)|>1]$. We refer to this as the singleton objective. To our knowledge, practical conformal prediction methods that aim to optimize the singleton objective have not yet been developed.
@@
    Empirical validation: We conduct detailed experiments on three image classification datasets (two versions of ImageNet and TissueMNIST) and LLM multiple-choice question answering. The results demonstrate that our method, which we call Singleton-Optimized Conformal Prediction (SOCOP), achieves a favorable balance between minimizing average set size and maximizing the frequency of singleton predictions compared to state-of-the-art baselines. Often, we can reduce the non-singleton probability by a large fraction (such as 20%) while only incurring a small increase in expected set size.
@@ 2 A Singleton-Optimized Nonconformity Score
The next and crucial observation is that the sets $S_{\eta,\gamma}$ from (2) are nested as a function of the Lagrange multiplier $\eta$, i.e., $S_{\eta_{1},\gamma}\subseteq S_{\eta_{2},\gamma}$ when $\eta_{1}<\eta_{2}$; see Lemma B.1 in Appendix B. This motivates us to define a nonconformity score via nested conformal prediction (Vovk et al., 2005; Gupta et al., 2022), where we aim to find the smallest $\eta$—and thus the smallest set $S_{\eta,\gamma}$—that contains the true label.
@@ 2.1 Geometric approach to computing the nonconformity score
Let the ordered sequence of vertices of the lower convex hull of $\mathcal{P}$ be $\{P_{v_{0}},P_{v_{1}},\dots,P_{v_{m}}\}$, where $v_{0}<v_{1}<\dots<v_{m}$ are indices from $\{0,\dots,K\}$. By construction, we have $v_{0}=0$ and $v_{m}=K$, since $\Gamma_{k}$ are strictly increasing and $g_{k}$ are non-decreasing with $k$. For $i=1,\dots,m$, define the slope of the edge connecting the vertices $P_{v_{i-1}}$ and $P_{v_{i}}$ as $\eta_{i}:=(g_{v_{i}}-g_{v_{i-1}})/(\Gamma_{v_{i}}-\Gamma_{v_{i-1}})$. To unify the analysis, we define $\eta_{0}:=0$ and $\eta_{m+1}:=+\infty$. The following theorem (with proof in Appendix B) characterizes the the jumps and slopes of $\kappa$. Figure 1 shows an example of a lower convex hull of a point set $\mathcal{P}$ for a probability vector with K=10K=10.555Red points indicate the hull vertices, and $\eta_{i}$ denote the corresponding slopes. The nonconformity scores are $r(x,y_{1})=\eta_{1}$, $r(x,y_{2})=\cdots=r(x,y_{7})=\eta_{2}$, $r(x,y_{8})=\eta_{3}$, $r(x,y_{9})=\eta_{4}$, and $r(x,y_{10})=\eta_{5}$.
@@
nonconformity scores and the final prediction sets. The form $r(x,y_{i})=\inf\{\eta\geqslant 0:y_{i}\in S_{\eta,\hat{p}(\cdot|x)}\}=\inf\{\eta\geqslant 0:\kappa(\eta;\hat{p}(\cdot|x))\geqslant i\}$ is equivalent to finding the smallest slope $\eta_{j}$ that leads to a prediction set of size $v_{j}\geqslant i$. To compute this, we can first find the vertices of the lower convex hull (which can be done with a standard approach, see Algorithm 2), and identify the correct slope; these can be performed in a single loop. Having an efficient algorithm to compute the nonconformity score is useful in a variety of conformal prediction methods, such as split conformal prediction (Papadopoulos et al., 2002), cross-conformal prediction (Vovk, 2015), Mondrian conformal prediction (Vovk et al., 2005), etc. In this paper, we will focus on split conformal prediction, which is one of the most popular and widely applicable methods.
@@ 3 Experiments
For our proposed SOCOP method, the hyperparameter $\lambda$ is selected by aiming to find a ”knee point” of the size-singleton probability curve on the tuning subset, as detailed in Section 3.1.1. The evaluation metrics we use are Coverage, Average Size, and P(size>>1).
@@ 3.1 Image classification on ImageNet
Evaluation on ImageNet-V2. We apply the same evaluation pipeline to ImageNet-V2 (Recht et al., 2019), which is a more challenging test dataset. Since this is a smaller dataset, we randomly sample three subsets of Imagenet-Val over 100 trials: one tuning subset of size 1K, one conformal calibration subset of size 4K, and one evaluation subset of size 4K. Table 10 in Appendix D.2 reports the results for all five models. The empirical findings are consistent with those on ImageNet-Val. Notably, the advantages of SOCOP are even more pronounced on this more challenging dataset. The variance of the coverage is higher due to having less data.
@@ 3.1.1 Effect of $\lambda$ and Hyperparameter Tuning
The right panel of Figure 2 summarizes this trade-off by plotting P(size>>1) against Average Size. In practice, one can choose $\lambda$ according to their own preference by drawing the tradeoff plot (the right panel of Figure 2) on their tuning dataset. For illustration, in our experiments from Section 3.1 we use the kneedle algorithm (Satopaa et al., 2011), which is a popular method for choosing points along a trade-off curve that come with favorable trade-offs. All five models exhibit the same pattern on both ImageNet-Val and ImageNet-V2, see Tables 5-15 in the Appendix.
@@ 3.4 Discussion
SOCOP reframes efficiency in conformal classification around the goal of producing singletons, deriving a nonconformity score from a geometric analysis of a Lagrangian relaxation of the singleton objective. This yields an $O(K)$ per-instance algorithm, enabling split conformal sets that preserve marginal coverage while substantially increasing singleton frequency with minimal impact on average size. Empirically, over image classification and LLM multiple-choice benchmarks, this reduces non-singleton rates significantly relative to length-optimized baselines at near-identical set sizes; suggesting that our method could be broadly useful in practice. In future work, it would be of interest to extend this method to more advanced conformal prediction methods, such as label-conditional or Mondrian conformal prediction (Vovk et al., 2005).
@@ Appendix B Auxiliary Results and Proofs
which contradicts with the optimality of $S_{\eta,\gamma}$. Therefore, $S_{\eta,\gamma}$ must be the set of top-$j$ labels for some $j$ that depends on $\eta$ and $\gamma$.
B.2 Lemma B.1
Lemma B.1 (Nested Sets Property).
For $\eta_{1}<\eta_{2}$, we have $S_{\eta_{1},\gamma}\subseteq S_{\eta_{2},\gamma}$.
Proof.
@@ D.1 ImageNet-Val
For this dataset, the effect of $\lambda$ on our SOCOP across all five models are reported in Table 5-9, respectively.
@@ D.2 ImageNet-V2
For this dataset, the effect of $\lambda$ on our SOCOP across all five models are reported in Table 11-15, respectively.

### v2

@@ 1 Introduction
However, average size is not necessarily the ideal measure of efficiency. Often, the most desirable outcome is an unambiguous prediction, a singleton set containing only one label. A set of size two or more may require additional human intervention or changing the workflow when used in downstream analysis, and thus brings an outsized cost. This motivates an alternative efficiency criterion, first conceptualized in Vovk et al. (2005) as the M-criterion, which seeks to minimize the probability of producing a non-singleton set, $\mathbb{P}_{X}[|C(X)|>1]$. We refer to this as the singleton objective.  [^1] Strictly speaking, a singleton set refers to a cardinality of exactly one ($|C(X)|=1$). In this work, we use the term “singleton objective” to broadly refer to the goal of minimizing the probability of returning multiple labels ($|C(X)|>1$). As we discuss below in our experiments, zero sets occur extremely rarely, and so the two objectives effectively coincide. To our knowledge, practical conformal prediction methods that aim to optimize the singleton objective have not yet been developed.
@@
    Empirical validation: We conduct detailed experiments on three image classification datasets (two versions of ImageNet and TissueMNIST) and LLM multiple-choice question answering. The results demonstrate that our method, which we call Singleton-Optimized Conformal Prediction (SOCOP), achieves a favorable balance between minimizing average set size and maximizing the frequency of singleton predictions compared to state-of-the-art baselines. Often, we can reduce the non-singleton probability by a large fraction (such as 20%) while only incurring a small increase in expected set size.  [^2] Code is available at https://github.com/TaoWangPenn/Singleton-Optimized-Conformal-Prediction.
@@ 2 A Singleton-Optimized Nonconformity Score
The next and crucial observation is that the sets $S_{\eta,\gamma}$ from (2) are nested as a function of the Lagrange multiplier $\eta$.
Lemma 2.2 (Nested Sets Property).
For $0\leqslant\eta_{1}<\eta_{2}$, we have $S_{\eta_{1},\gamma}\subseteq S_{\eta_{2},\gamma}$.
This motivates us to define a nonconformity score via nested conformal prediction (Vovk et al., 2005; Gupta et al., 2022), where we aim to find the smallest $\eta$—and thus the smallest set $S_{\eta,\gamma}$—that contains the true label.
@@ 2.1 Geometric approach to computing the nonconformity score
Let the ordered sequence of vertices of the lower convex hull of $\mathcal{P}$ be $\{P_{v_{0}},P_{v_{1}},\dots,P_{v_{m}}\}$, where $v_{0}<v_{1}<\dots<v_{m}$ are indices from $\{0,\dots,K\}$. By construction, we have $v_{0}=0$ and $v_{m}=K$, since $\Gamma_{k}$ are strictly increasing and $g_{k}$ are non-decreasing with $k$. For $i=1,\dots,m$, define the slope of the edge connecting the vertices $P_{v_{i-1}}$ and $P_{v_{i}}$ as $\eta_{i}:=(g_{v_{i}}-g_{v_{i-1}})/(\Gamma_{v_{i}}-\Gamma_{v_{i-1}})$. To unify the analysis, we define $\eta_{0}:=0$ and $\eta_{m+1}:=+\infty$. The following theorem (with proof in Appendix B) characterizes the the jumps and slopes of $\kappa$. Figure 1 shows an example of a lower convex hull of a point set $\mathcal{P}$ for a probability vector with K=10K=10.777Red points indicate the hull vertices, and $\eta_{i}$ denote the corresponding slopes. The nonconformity scores are $r(x,y_{1})=\eta_{1}$, $r(x,y_{2})=\cdots=r(x,y_{7})=\eta_{2}$, $r(x,y_{8})=\eta_{3}$, $r(x,y_{9})=\eta_{4}$, and $r(x,y_{10})=\eta_{5}$.
@@
nonconformity scores and the final prediction sets. The form $r(x,y_{i})=\inf\{\eta\geqslant 0:y_{i}\in S_{\eta,\hat{p}(\cdot|x)}\}=\inf\{\eta\geqslant 0:\kappa(\eta;\hat{p}(\cdot|x))\geqslant i\}$ is equivalent to finding the smallest slope $\eta_{j}$ that leads to a prediction set of size $v_{j}\geqslant i$. Intuitively, the slope $\eta$ represents the “price” per unit of coverage relative to the set-size penalty. Navigating the lower convex hull corresponds to finding the minimum price required to “purchase” enough coverage to include the target label $y_{i}$ in the set. To compute this, we can first find the vertices of the lower convex hull (which can be done with a standard approach, see Algorithm 2), and identify the correct slope; these can be performed in a single loop. Having an efficient algorithm to compute the nonconformity score is useful in a variety of conformal prediction methods, such as split conformal prediction (Papadopoulos et al., 2002), cross-conformal prediction (Vovk, 2015), Mondrian conformal prediction (Vovk et al., 2005), etc. In this paper, we will focus on split conformal prediction, which is one of the most popular and widely applicable methods.
@@ 3 Experiments
For our proposed SOCOP method, the hyperparameter $\lambda$ is selected by aiming to find a ”knee point” of the size-singleton probability curve on the tuning subset, as detailed in Section 3.1.1. The evaluation metrics we use are Coverage, Average Size, and P(size>>1).  [^10] We also evaluated the empty set rate, $P(|C(X)|=0)$, across all experiments. We observed that empty sets occur in fewer than 0.01% of test instances for all methods, with the exception of RAPS (where they can be slightly larger but still insignificant, reaching $\approx 0.1\%$). Consequently, we omit this metric from the results as its impact is negligible.
@@ 3.1 Image classification on ImageNet
Evaluation on ImageNet-V2. We apply the same evaluation pipeline to ImageNet-V2 (Recht et al., 2019), which is a more challenging test dataset. This dataset was constructed by re-collecting images with a new sampling pipeline, introducing a natural distribution shift that typically results in a significant drop in accuracy for models trained on the original ImageNet dataset. Since this is a smaller dataset, we randomly sample three subsets of Imagenet-Val over 100 trials: one tuning subset of size 1K, one conformal calibration subset of size 4K, and one evaluation subset of size 4K. Table 11 in Appendix D.2 reports the results for all five models. The empirical findings are consistent with those on ImageNet-Val. Notably, the advantages of SOCOP are even more pronounced on this more challenging dataset. The variance of the coverage is higher due to having less data.
@@ 3.1.1 Effect of $\lambda$ and Hyperparameter Tuning
The right panel of Figure 2 summarizes this trade-off by plotting P(size>>1) against Average Size. In practice, one can choose $\lambda$ according to their own preference by drawing the tradeoff plot (the right panel of Figure 2) on their tuning dataset. For illustration, in our experiments from Section 3.1 we use the kneedle algorithm (Satopaa et al., 2011), which is a popular method for choosing points along a trade-off curve that come with favorable trade-offs. All five models exhibit the same pattern on both ImageNet-Val and ImageNet-V2, see Tables 6-16 in the Appendix.
@@
3.1.2 Adaptiveness on ImageNet
In this experiment, we evaluate the size-stratified coverage violation (SSCV) introduced by Angelopoulos et al. (2021) as a measure of adaptiveness and conditional coverage violation. Following Angelopoulos et al. (2021), we adopt the same set-size strata : 0-1, 2-3, 4- 10, 11-100, and 101-1000, and to maximize adaptiveness, we choose the hyperparameter $\lambda$ to minimize SSCV on the tuning set for RAPS and SOCOP. The details of the hyperparameter grid are provided in Appendix E. The results are reported in Table 2.
Table 2: Evaluation results for the SSCV metric on ImageNet-Val, with the same protocol as in Table 1.
Method Coverage Avg Size $P(\textnormal{size}>1)$ SSCV
ResNet152-v2
Plug-In $0.969\pm 0.003$ $47.362\pm 7.138$ $469\pm 0.025 0.0$$046\pm 0.001$
RAPS $0.950\pm 0.002$ $8.568\pm 1.580$ $448\pm 0.012 0.0$$031\pm 0.011$
Pure Singleton $0.950\pm 0.002$ $250.539\pm 4.554$ $250\pm 0.005 0.0$$050\pm 0.000$
Least Ambiguous Sets $0.950\pm 0.002$ $2.279\pm 0.046$ $0.467\pm 0.007$ $0.197\pm 0.026$
SOCOP (ours) $0.950\pm 0.002$ $3.372\pm 0.198$ $304\pm 0.008 0.0$$039\pm 0.009$
ViT-h-14
Plug-In $0.976\pm 0.001$ $8.529\pm 0.805$ $356\pm 0.008 0.0$$048\pm 0.002$
RAPS $0.950\pm 0.002$ $7.652\pm 2.259$ $319\pm 0.007 0.0$$047\pm 0.003$
Pure Singleton $0.950\pm 0.003$ $136.219\pm 4.980$ $135\pm 0.005 0.0$$050\pm 0.000$
Least Ambiguous Sets $0.950\pm 0.002$ $1.291\pm 0.011$ $0.224\pm 0.006$ $0.126\pm 0.119$
SOCOP (ours) $0.950\pm 0.002$ $1.519\pm 0.068$ $155\pm 0.006 0.0$$041\pm 0.016$
Our method SOCOP and RAPS achieve the smallest SSCV among the methods compared. However, the average size of RAPS increases drastically (from $\approx 3.2$ to $\approx 8.6$ for Resnet152-v2 and from $\approx 1.4$ to $\approx 7.7$ for ViT-h-14), while our SOCOP method maintains a reasonably small average size and a significantly lower non-singleton probability, demonstrating that SOCOP can achieve adaptivity without sacrificing efficiency.
@@ 3.4 Discussion
SOCOP reframes efficiency in conformal classification around the goal of producing singletons, deriving a nonconformity score from a geometric analysis of a Lagrangian relaxation of the singleton objective. This yields an $O(K)$ per-instance algorithm, enabling split conformal sets that preserve marginal coverage while substantially increasing singleton frequency with minimal impact on average size. Empirically, over image classification and LLM multiple-choice benchmarks, this reduces non-singleton rates significantly relative to length-optimized baselines at near-identical set sizes; suggesting that our method could be broadly useful in practice.
In future work, it would be of interest to extend this method to more advanced conformal prediction methods, such as label-conditional or Mondrian conformal prediction (Vovk et al., 2005). Furthermore, a challenging but interesting theoretical direction is to design a nonconformity score intrinsically targeted for conditional coverage. This would likely entail retracing the derivation of SOCOP starting from a conditional-aware optimization objective, such as the one from Gibbs et al. (2025). Finally, while our current protocol uses a separate tuning set to maintain validity, future work could investigate data-dependent selection of $\lambda$ using the calibration set directly to improve data efficiency, while accounting for the resulting tuning bias (Zeng et al., 2025).
@@ Appendix A Additional related work
Recent frameworks [Liang et al., 2024, Yang and Kuchibhotla, 2025] optimize efficiency by selecting a nonconformity score from a pre-specified candidate set that minimizes a target loss. This differs fundamentally from our approach, which uses the loss to derive the score directly. Applying our composite loss within these selection frameworks faces two practical difficulties: (1) the resulting performance is strictly bounded by the pre-defined candidate pool; and (2) this would introduce the additional difficulty of selecting the hyperparameter $\lambda$, and their framework would have to be potentially extended to allow the selection of not just non-conformity scores but also loss functions.
@@ Appendix B Auxiliary Results and Proofs
which contradicts with the optimality of $S_{\eta,\gamma}$. Therefore, $S_{\eta,\gamma}$ must be the set of top-$j$ labels for some $j$ that depends on $\eta$ and $\gamma$. ∎
B.2 Proof of Lemma 2.2
@@ D.1 ImageNet-Val
For this dataset, the effect of $\lambda$ on our SOCOP across all five models are reported in Table 6-10, respectively.
@@ D.2 ImageNet-V2
For this dataset, the effect of $\lambda$ on our SOCOP across all five models are reported in Table 12-16, respectively.
@@ D.3 CPL Method
Appendix E Hyperparameter Grid
For RAPS, we follow Angelopoulos et al. [2021], using the gird $\lambda\in\{0.001,0.01,0.1,0.2,0.5\}$ to optimize set size and a grid with smaller values $\lambda\in\{0.00001,0.0001,0.0008,0.001,0.0015,0.002\}$ to optimize SSCV. For our SOCOP method, we use a linearly spaced grid of 15 values over over [0.05, 1.0] to optimize the balance between set size and non-singleton rate, and a linearly spaced grid of 15 values over $[0.005,0.1]$ to optimize SSCV.